OpenStudy (anonymous):

y = 3x - 4 x = 9 A. (9, 23) B. (1, -1) C. (0, -4) D. (9, 31). How do I solve this?

OpenStudy (nfcfox):

Let's read the problem. Ok so they give us y= 3x-4 and x= 9. What we want to do is use direct substitution. The best way of doing this would to just take the 9 from x= 9 and plug it right into our y= equation (we want to find the y coordinate in order to get the ordered pair). Using direct substation we will get: y= 3(9)-4 Simplify! y= 27-4 y= 23 Now that we have our y we are done! Just take the x that we already know and make the ordered pair. (9,23) So the answer is is A.